OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 10, x = 3. Find y when x = 5

OpenStudy (anonymous):

"y varies inversely with twice the value of x." \[y=k\frac{ 1 }{ 2x }\]

OpenStudy (anonymous):

When y = 10, x = 3. \[(10)=k\frac{ 1 }{ 2(3) }\]\[10=k\frac{ 1 }{ 6 }\]both sides times 6:\[60=k\]

OpenStudy (anonymous):

10=k(1/6) 60=K Y=60(1/5) Y=12

OpenStudy (anonymous):

thanks up so much dss